OpenStudy (anonymous):

Given that y = 4kx^2 - 6x + (k-4), find the values of k, for which the equation y = 0 has real roots. So i did b^2 - 4ac>0 and i got (1 + 2k)(9 - 2k)>0 Now what??

OpenStudy (campbell_st):

you have factorised correctly so look at k ignoring the inequality the solutions are k = -1/2 and 9/2 I would pick an value between the k values to test say k = 0 then the quadratic (1 + 2k)(9 - 2k)> 0 when k = 0 so the inequality is \[-\frac{1}{2} < k < \frac{9}{2}\]

OpenStudy (anonymous):

YUP i got that :)